The efficiency of a machine is 60%, calculate the workdone by a man using this machine to raise a load of 250kg through a vertical distance of 2.5m(g=10ms-2)
1
Expert's answer
2020-01-13T11:02:35-0500
"\rm The\: efficiency=\frac{Output\:work}{Input\:work}"

"\eta=\frac{mgh}{W}"

Hence, the work done by a man


"W=\frac{mgh}{\eta}=\frac{250\times 10\times 2.5}{0.6}=10417\:\rm J"

Answer: "10417\:\rm J"